OpenStudy (anonymous):

Might anyone be able to help, please. How are operations with radical expressions similar to operations with variables?

OpenStudy (turingtest):

well rules tend to be rather similar\[a\cdot b=b\cdot a\\\sqrt a\cdot\sqrt b=\sqrt b\cdot\sqrt a\](this is called the commutative property)